OpenStudy (anonymous):

3x + 5y = 25 2x − 7y = −4

OpenStudy (saifoo.khan):

Where are u stuck in that?

hero (hero):

y = (25 - 3x)/5 y = (4+2x)/7 Set y = y: (25 - 3x)/5 = (4+2x)/7 7(25-3x) = 5(4+2x) 175 - 21x = 20 + 10x 155 = 31x x = 155/31 y = something

OpenStudy (anonymous):

Its the addition method and I was absent and my teacher told me to try to figure it out, please help